OpenStudy (anonymous):

Solve 4(2x1-)=3(x-3)-7

OpenStudy (anonymous):

3(x-3)-7

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) to solve for x for the equation 4(2x-1) = 3(x-3) - 7 first distribute so... 8x - 4 = 3x - 9 - 7 then add like terms 8x - 4 = 3x - 16 then subtract 3x from both sides -3x -3x 5x - 4 = -16 then add 4 to both sides +4 +4 5x = -12 then divide both sides by 5 /5 /5